OpenStudy (anonymous):

is cos²x/2 = 1+cosx/2 AND cosx/2 = ±√1+cosx/2 Same thing??

jimthompson5910 (jim_thompson5910):

If you mean \[\Large \cos^2\left(\frac{x}{2}\right) = 1+\cos\left(\frac{x}{2}\right)\] is the same as \[\Large \cos\left(\frac{x}{2}\right) = \pm\sqrt{1+\cos\left(\frac{x}{2}\right)}\] then you are correct

OpenStudy (anonymous):

OKay thank you so much!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

is it possible for you to check my work?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

so it said to verify the identity of 2cscx cos²(x/2)= sinx/1-cosx

OpenStudy (anonymous):

i will write down the steps i know so far

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

somy first step i did was this!

OpenStudy (anonymous):

but i dont know what is next....could you please help me

jimthompson5910 (jim_thompson5910):

From there, multiply top and bottom of (1+cos)/(sin) by (1-cos) Tell me what you get

OpenStudy (anonymous):

okay i am almost done please wait for a moment!

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

you nailed it, very nice work

jimthompson5910 (jim_thompson5910):

I'm sure you already know this, but it's important to keep in mind. When dealing with identities, you only manipulate one side. So you convert one side to the other.

OpenStudy (anonymous):

okay! thank you so much!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Sorry to ask you again but How do you verify the identities of cos²(x/2) = secx+1 / 2secx ??? i got steps done if you could check

OpenStudy (anonymous):

How do you verify the identities of cos²(x/2) = secx+1 / 2secx ???

jimthompson5910 (jim_thompson5910):

sec = 1/cos

jimthompson5910 (jim_thompson5910):

so cos*sec = cos*(1/cos) = cos/cos = 1

jimthompson5910 (jim_thompson5910):

this means (1+cos)*sec = sec + 1

OpenStudy (anonymous):

sorry i am really lost

jimthompson5910 (jim_thompson5910):

do you see how sec = 1/cos

OpenStudy (anonymous):

yes i get that part

jimthompson5910 (jim_thompson5910):

do you see how cos*sec = cos*(1/cos) = cos/cos = 1 ?

OpenStudy (anonymous):

but before we started the this question dont you multiply the conjugate? so wouldnt it be 1+cosx (1-cosx) and 2(1-cosx )???

jimthompson5910 (jim_thompson5910):

you could do that (it's perfectly valid), but you'll be getting further away from what we want (which has a sec in it)

jimthompson5910 (jim_thompson5910):

They multiplied by sec/sec because they wanted to introduce a 'sec' term

OpenStudy (anonymous):

ok...could you explain a bit in detail about the sec term? I am so sorrrrry :((

jimthompson5910 (jim_thompson5910):

we have (1+cos)/2

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

we want a sec term in the denominator (along with the 2), so we multiply 2 by sec

jimthompson5910 (jim_thompson5910):

but we must do the same to the numerator to keep things balanced (and equivalent)

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

so do you see how/why they got the answer?

OpenStudy (anonymous):

hmm,,? like this ?

jimthompson5910 (jim_thompson5910):

close, distribute that "sec" through to every term in the numerator

jimthompson5910 (jim_thompson5910):

I'm sure you know this (by now), but when I say "sec" I mean "sec(x)", but I find it easier to type "sec"

OpenStudy (anonymous):

okay im kind of lost

jimthompson5910 (jim_thompson5910):